OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (anonymous):

3(x+1) = -2(x-1) + 6 3x+ 3 = -2x + 2 +6 3x+3= -2x + 8 3x + 2x = -3 + 8 5x = 5 x= 5/5 x= 1 :)

OpenStudy (anonymous):

3(x + 1) = -2(x - 1) + 6 Distributing... 3x + 3 = -2x + 2 + 6 3x + 3 = -2x + 8 Try it from here. If you are still having problems, Ill be happy to help

OpenStudy (anonymous):

@rani Please do not give out the answer, but help the OP in answering it by themselves

OpenStudy (anonymous):

yeah, your right... they need to try it too, sorry about that ;)

OpenStudy (anonymous):

@chandhooguy @rani i'm not looking for the answer just for some help! but thanks guys!!

OpenStudy (anonymous):

@rani It's ok, it is a common mistake :)